Select the vibrations that should be infrared active.
Answer #1
Concepts and reason

Vibration spectrum is the spectrum occursfrom the transition induced between the vibrationalenergy levels of a molecule. These spectra occur in the spectral range of 4000-500 cm-1. Compounds that have permanent dipole moment are IR active.

Fundamentals

Main criteria for molecules to show an infrared spectrum is that the dipole moment of the molecule must change during vibration.Change in dipole occurs mostly in asymmetric molecules.

Molecules that possess permanent dipole moment are IR inactive.

CH3CH2CCCH2CH3{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{C}} \equiv {\rm{CC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{3}}}

HỌC
––––CH,

This is a symmetric molecule, so this molecule shows no dipole moment.

(CH3)2C=O{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{2}}}{\rm{C = O\;\;}} - CH3
H3C—
ć=
0

This is aasymmetric compound, so it has dipole moment.

trans4octene{\rm{trans - 4 - octene}}\, - trans-4-octane

This compound is symmetric, so it shows zero dipole moment.

(CH3CH2)3CCI{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}} \right)_{\rm{3}}}{\rm{C -- CI\;\;\;}} - CH2CH3
H3CH2C—
¢-
C1
CH2CH3

This compound is asymmetric, so it shows dipole moment.

CH3CH2CH2C=CH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{C = CH\;\;\;\;}} - H3CH2CH2C—C=
CH

This compound has dipole moment because it is not symmetric.

IR active compounds

IR inactive compounds

(CH3)2C=O(C=Ostretch){\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{2}}}{\rm{C = O\;\;}}\left( {{\rm{C = O stretch}}} \right)

CH3CH2CCCH2CH3(CCstretch){\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{C}} \equiv {\rm{CC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{(C}} \equiv {\rm{C stretch)}}

(CH3CH2)3CCI(CClstretch){\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}} \right)_{\rm{3}}}{\rm{C -- CI\;\;\;}}\left( {{\rm{C - Cl stretch}}} \right)

trans4octene(C=Cstretch){\rm{trans - 4 - octene}}\,{\rm{(C = C}}\,{\rm{stretch)}}

CH3CH2CH2C=CH(C=Cstretch){\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{C = CH\;\;\;\;}}\left( {{\rm{C = C stretch}}} \right)

Ans:

OCH,CH.C=CCH.CH,C=C stretch)
ECH ), C = 0 (C= 0 stretch)
trans-4-octene(C=Cstretch)
D(CH, CH), C-CI (C-C1 stretch)
CH.CH,CH,C